Please help me to answer this problem.
There is a bowl of candy sitting on our kitchen table. This morning Nadia takes one-sixth of the candy. Later that morning Peter takes one-fourth of the candy that’s left. This afternoon, Andrew takes one-fifth of what’s left in the bowl and finally Anne takes one-third of what is left in the bowl. If there are 16 candies left in the bowl at the end of the day, how much candy was there at the beginning of the day?

Answer by htmentor(1343)

Let C = the original number of candies
The remaining number of candies after all the people have taken some is 16.
After Nadia takes 1/6 there are 5C/6 left. After Peter takes 1/4 of the
remaining number, there are (3/4)(5/6)C, and so on...
Thus R = (5/6)(3/4)(4/5)(2/3)C -> R = C/3 = 16
Thus C = 48
Check: Nadia: 40 left, Peter: 30 left, Andrew: 24 left, Anne: 16